OpenStudy (anonymous):

For the function, f(x) = -log(3-x)+4 a) State the domain and range: b) Solve for the x and y intercepts: c) List any asymptotes: d) Sketch and include properties in your sketch: PLEASE SHOW YOUR WORK :)

OpenStudy (across):

Do you know what the domain of the log function is?

OpenStudy (anonymous):

I have not a clue on how to do this

OpenStudy (across):

Okay, first, it'd be good to know that the log function is defined on the domain (0,∞).

OpenStudy (across):

That means that, no matter what, the 3-x inside the log function cannot be neither zero nor negative. In other words, 3-x>0. Can you solve for x?

OpenStudy (anonymous):

I'm confused

OpenStudy (across):

Solve for x\[3-x>0\]

OpenStudy (anonymous):

I haven't learned that yet, I don't think

OpenStudy (across):

You have:\[3-x>0.\]Add x to both sides:\[3-x+x>0+x,\]\[3-\not{x}+\not{x}>x,\]\[3>x.\]

OpenStudy (across):

If x < 3, do you know what the domain of the function f(x)=-log(3-x)+4 is?

OpenStudy (anonymous):

so the domain and range would be...?

OpenStudy (anonymous):

x<3?

OpenStudy (across):

That means that x cannot be greater than three. In other words, your domain is (-∞,3) or x<3, that's correct.

OpenStudy (anonymous):

so i would write it as xer, x<3?

OpenStudy (across):

Now, the range is a totally different story. For it, you need to calculate how the function behaves as x→3 and x→-∞.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would you do that?

OpenStudy (across):

Well, you can evaluate that\[\lim_{x\to3}-\log(3-x)+4=\infty.\]How about as x→-∞?

OpenStudy (anonymous):

I'm confused once again.

OpenStudy (across):

\[\lim_{x\to-\infty}-\log(3-x)+4=?\]

OpenStudy (anonymous):

what's lim?

OpenStudy (across):

It's the limit of a function as its independent variable(s) approach a certain/certain number(s).

OpenStudy (anonymous):

ok so the range is that y cannot be negative?

OpenStudy (anonymous):

I'm so sorry but I have to eat dinner, can you answer all of these questions? I'll be back ASAP.

OpenStudy (across):

Well, inspecting\[\lim_{x\to-\infty}-\log(3-x)+4,\]we'll see that 3-x is going to be really huge, log(3-x) is going to be really huge, -log(3-x) is going to be really small, and adding 4 to it won't do anything significant. Thus\[\lim_{x\to-\infty}-\log(3-x)+4=-\infty.\]Since we obtained that the range fits in between −∞ and ∞, then we can conclude that the range of the function is (−∞,∞) or the set of all real numbers\[\mathbb{R}.\]

OpenStudy (across):

I'm a little confused: If you couldn't solve a very, very trivial inequality, namely 3 - x > 0, then why are you working problems which require you know what a limit is?

OpenStudy (across):

I'm going to give you a few hints. To find the y-intercept(s), you simplify the following equation:\[y=-\log(3)+4.\]To find the x-intercept(s), you solve the following equation for x:\[0=-\log(3-x)+4.\]Since we found the range of the function to go from -∞ to ∞, then there are no horizontal asymptotes. However, at which point is the function undefined? This point is a vertical asymptote. Upon knowing all these things, sketching the graph should be a piece of cake. Have fun!

OpenStudy (anonymous):

helloo?